Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day14


Crew #108 Log Book for December 16, 2011

Photos of the day





Commander's Check-In Report

Date and time:

16 December 2011

Written by: 

Charlotte Poupon, Crew 108, Commander

Crew Physical Status:

5 all right

Time Departed/Returned from EVA:

EVA # 24 / from 09:55 to 11:10


Brief Narrative of Field Mission Results:

Not applicable

EVA Data/Interpretations:

Not applicable

Engineering/Hab Maintenance:

Engineers to rebuild a new breathing system. It is pretty efficient.

Report Transmission Schedule:

All reports by 22:00.

Plans for Tomorrow:

Handover to Crew 109

Inventory:

Not applicable

Support Requested:

Not applicable

Miscellaneous:

Commander's Report

Date and time:

16 December 2011

Written by:

Charlotte Poupon, Crew 108, Commander

Text:

Today we get the Hab ready for the rotation 109.

On the morning, we've cleaned the kitchen and we've organized food, following Jean Hunter's instruction.

The 2 empty bins were put in the loft, while a third ones stands for the leftover.

Floor was swept.

Lab was cleaned as well.

Thought of the day :

I heard the mission bell, and I was thinking to myself, this could be heaven or this could be hell…

It was piece of heaven here.

Science Report

Date and Time:

16 December 2011, 14:41

Written by: 

Usha Lingappa, Crew 108, Astrobiologist

To prepare for the end of the mission tomorrow, I cleaned up the lab and sterilized all my cultures.

The limited amount of reagents and equipment in the Hab precluded the application of molecular biology techniques that would have been my next step in studying the area's microbiology. While molecular techniques provide us with more information on environmental microbiology than any other techniques at our disposal, they are only relevant to life as we know it (organisms with genetic material encoded in nucleic acids). It's very possible that life on Mars could be completely different. If I have the opportunity to continue simulated Martian microbiological investigations in the future, I would rather focus on alternative techniques for identifying microbial life including selective cultivation methods and chemical analyses looking for evidence of metabolic activity.


For future crews- incubation temperature range of lab oven:

Internal temperature when off: ~15ºC

Minimum heat: ~35ºC

Every mark = roughly 2ºC

Maximum temperature: ~200ºC


The oven temperature control had no values, so I took the liberty of writing these numbers on pieces of tape that I stuck on the oven.

Science Report

Date and Time: 

15 December 2011, 14:41

Written by: 

Usha Lingappa, Crew 108, Astrobiologist

To prepare for the end of the mission tomorrow, I cleaned up the lab and  sterilized all my cultures.

The limited amount of reagents and equipment in the Hab precluded the application of molecular biology techniques that would have been my next step in studying the area's microbiology. While molecular techniques provide us with more information on environmental microbiology than any other techniques at our disposal, they are only relevant to life as we know it (organisms with genetic material encoded in nucleic acids). It's very possible that life on Mars could be completely different. If I have the opportunity to continue simulated Martian microbiological investigations in the future, I would rather focus on alternative techniques for identifying microbial life including selective cultivation methods and chemical analyses looking for evidence of metabolic activity.

For future crews- incubation temperature range of lab oven:

Internal temperature when off: ~15ºC

Minimum heat: ~35ºC

Every mark = roughly 2ºC

Maximum temperature: ~200ºC


The oven temperature control had no values, so I took the liberty of writing these numbers on pieces of tape that I stuck on the oven.

Engineering Report

Date and time: 

16 December 2011, 19:45 

Written by: 

Mike Lotto, Crew 108, Engineer

Generator/Electricity:
Did you have any electricity outage? (Yes/No): No

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): 0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5
Notes/Comments:

Oil:
ATV Oil Quantity (# of Quarts in storage): 1.25
Car Oil Quantity (# of Quarts in storage): 1.25

Water Status:
Outside Potable Water Tank Level (inches from bottom): 10
Trailer Potable Water Tank Level (inches from bottom):
Approximately 30 inches (or 70% full)
Hab Potable Water Tank Level (inches from bottom): 9
Potable Water Meter Reading: Unknown
In to GreenHab Meter Reading: Offline
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: Offline
Condition of plants in Tank 2: Offline
Condition of plants in Tank 3: Offline
Greenhab Notes/Comments: GreenHab is offline this season.

Observatory:
Telescope Used? (Yes/No): No
Observatory Notes/Comments:

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): Yes
Oil Checked (yes/no): No
Tire Status: Good
Odometer at end of day: 125475 miles
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: Yes
Fuel Consumed: 0.1 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 13:59
Comments:

ATV 2 (Spirit):
Used: Yes
Oil Checked: Yes
Fuel Consumed: 0.1 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 09:28
Comments:

ATV 3 (Viking I):
Used: Yes
Oil Checked: Yes
Fuel Consumed: 0.1 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 10:44
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 70
Actual temperature upstairs (Farenheit): 69

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 57

Computers/Networking Infrastructure:
Notes/comments on internet/computers?: None

General Engineering Notes/Comments: Good progress made on EVA suit 4 repairs.

Journalist Report

Date and time: 

16 December 2011, 23:59

Written by: 

Ashley Dale, Crew 108, Engineer

This morning, the boys went out to play with rocks on Olympus Mons and the girls cleaned up the Hab. Typical! Lotto and I worked on completing the construction of a new EVA backpack after one of the crew slipped on ice and shattered a previous one. The evening was spent being terrified by the film Alien. We also had one final stint at being awed by the stars. With such a cold and clear night, it really was quite a light-show – a great send-off. Today has been the final day Crew 108 is on Mars. Being fortunate enough to be privileged with living our share at the MDRS, we simply have no right to be gloomy about leaving simulation. But some strong bonds have been made, and it will be difficult to part ways for now. Each of us has learned some valuable lessons about ourselves. It has been a truly wonderful, unique, and character-building experience. These are memories I will cherish.

Crew 108, signing out.


Chef's Report

Date:

16 December 2011 

Written by:

Ashley Dale, Crew 108, Engineer

1a) Today was a non-cooking day.

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

1/2 tin of Long Grain and Wild Rice Pilaf
1 tin of pre-cooked Pink Salmon
1 tin of pre-cooked Oysters
6 x teaspoon of freeze-dried cheddar cheese powder

3) List any main meal foods not finished by the crew
1/3 of a bowl not finished.
1 portion leftover in cooking pot - stored in fridge for later.

6) Comments/questions for Kim and Jean, the food study investigators

This meal was very quick and easy to do.